# InstructIR: High-Quality Image Restoration Following Human Instructions

### v4

@@ 5 Multi-Task Ablation Study
Table 6: Summary ablation study on the multi-task variants of InstructIR that tackle from 3 to 7 tasks. We report PSNR/SSIM.
@@
How does 6D work?
Besides the 5 basic tasks -as previous works-, we include single image super-resolution (SISR). For this, we include as training data the DIV2K [1]. Since our model does not perform upsampling, we use the Bicubic degradation model [1, 15] for generating the low-resolution images (LR), and the upsampled versions (HR) that are fed into our model to enhance them. Adding this extra task increases the performance on deblurring –a related degradation–, without harming notably the performance on the other tasks. However, the performance on SR benchmarks is far from classical super-resolution methods [1, 44].
How does 7D work?
Finally, if we add image enhancement –a task not related to the previous ones i.e. inverse problems– the performance on the restoration tasks decays slightly. However, the model still achieves state-of-the-art results. Moreover, as we show in Table 5, the performance on this task using the MIT5K [5] Dataset is notable, while keeping the performance on the other tasks. We achieve similar performance to classical task-specific methods.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Clean up my image, “Get rid of the grain “Remove the strange spots
it’s too fuzzy.” in my photo” on my photo”
Refer to caption Refer to caption Refer to caption Refer to caption
“Retouch this image and “Reduce the motion “Please get rid of “Reduce the fog in
improve colors” in this shot” the raindrops” this landmark”
Figure 6: Instruction-based Image Restoration. InstructIR understands a wide a range of instructions for a given task (first row). Given an adversarial instruction (second row), the model performs an identity –we did not enforce this during training–. Images from BSD68 [52].
Refer to caption Refer to caption Refer to caption
Input (1)“Clear the rain from my picture” $\longrightarrow$ (2)“Make this photo look breathtaking”
Refer to caption Refer to caption Refer to caption
(1) “Retouch it as a photographer” $\longrightarrow$ (2) “Can you remove the raindrops?” $\longrightarrow$ (3) “Increase the resolution and details”
Refer to caption Refer to caption Refer to caption
Input (1)“My image is too dark, can you fix it?” (2)“Apply tone-mapping to the photo”
Figure 7: Multiple Real Instructions. We can prompt multiple instructions (in sequence) to restore and enhance the images. This provides additional control. We show two examples of multiple instructions applied to the “Input” image -from left to right-.
@@
We compare with task-specific methods for image enhancement in Table 5, and for low-light in image enhancement in 7. We provide extensive comparisons for image denoising in Table 8. Also, in Table 9 we show comparisons with classical methods for deblurring and dehazing. Our multi-task method is better than most task-specific methods, yet it is still not better than SOTA.
6 On the Effectiveness of Instructions
Thanks to our integration of human instructions, users can control how to enhance the images. We show an example in Figure 5, where the input image has three different degradations, and we aim to focus on a particular one. Although these results do not offer great reconstruction, we believe it is a promising direction that illustrates the effectiveness of instruction guidance for image restoration and enhancement. We provide more results in Figures 6 and 7, where we show the potential of our method to restore and enhance images in a controllable manner.
Table 7: Quantitative comparisons with state-of-the-art methods on the LOL dataset [83] (low-light enhancement). Table based on [81].
@@ 6 On the Effectiveness of Instructions
This implies an advancement w.r.t classical (deterministic) image restoration methods. Classical deep restoration methods lead to a unique result, thus, they do not allow to control how the image is processed. We also compare InstructIR with InstructPix2Pix [4] in Figure 8.
Qualitative Results.
We provide diverse qualitative results for several tasks. In Figure 9, we show results on the LOL [83] dataset. In Figure 10, we compare methods on the motion deblurring task using the GoPro [57] dataset. In Figure 11, we compare with different methods for the dehazing task on SOTS (outdoor) [41]. In Figure 12, we compare with image restoration methods for deraining on Rain100L [21]. Finally, we show denoising results in Figure 13. In this qualitative analysis, we use our single InstructIR-5D model to restore all the images.
@@
Discussion on Instruction-based Restoration
In Figure 8 we compare with InstructPix2Pix [4]. Our method is notably superior in terms of efficiency, fidelity and quality. We can conclude that diffusion-based methods [67, 53, 4] for image manipulation require complex “tuning” of several (hyper-)parameters, and strong regularization to enforce fidelity and reduce hallucinations. InstructPix2Pix [4] cannot solve inverse problems directly –although it has a good prior for solving Inpainting–, which indicates that such model require restoration-specific training (or fine-tuning).
Limitations
Our method achieves state-of-the-art results in five tasks, proving the potential of using instructions to guide deep blind restoration models. However, we acknowledge certain limitations. First, in comparison to diffusion-based restoration methods, our current approach would not produce better results attending to perceptual quality. Second, our model struggles to process images with more than one degradation (i.e. real-world images), yet this is a common limitation among the related restoration methods. Third, as previous all-in-one methods, our model only works with in-distribution degradations, thus it will not work on unseen artifacts. Nevertheless, these limitations can be surpassed with more realistic training data.
7 Conclusion
We present the first approach that uses human-written instructions to guide the image restoration models. Given natural language prompts, our model can recover high-quality images from their degraded counterparts, considering multiple degradation types. InstructIR achieves state-of-the-art results on several restoration tasks, demonstrating the power of instruction guidance. These results represent a novel benchmark for text-guided image restoration.n
Acknowledgments This work was partly supported by the The Humboldt Foundation (AvH). Marcos Conde is also supported by Sony Interactive Entertainment, FTG.
Instruction: “Reduce the noise in this photo” – Basic & Precise
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Remove the tiny dots in this image” – Basic & Ambiguous
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Improve the quality of this image” – Real user (Ambiguous)
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “restore this photo, add details” – Real user (Precise)
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Enhance this photo like a photographer” – Basic & Precise
Refer to caption Refer to caption Refer to caption Refer to caption
Input InstructIR (ours) InstructPix2Pix  $S_{I}\!=\!5$ InstructPix2Pix  $S_{I}\!=\!7$
Figure 8: Comparison with InstructPix2Pix [4] for instruction-based restoration using the prompt. Images from the RealSRSet [80, 44]. We use our 7D variant. We run InstructPix2Pix [4] using two configurations where we vary the weight of the image component hoping to improve fidelity: $S_{I}\!=\!5$ and $S_{I}\!=\!7$ (also known as Image CFG), this parameters helps to enforce fidelity and reduce hallucinations.
Figure 9: Low-light Image Enhancement Results. We compare with other methods on LOL [83] (748.png).
Figure 10: Image Deblurring Results. Comparison with other methods on the GoPro [57] dataset (GOPR0854-11-00-000001.png).
Figure 11: Image Dehazing Results. Comparison with other methods on SOTS [41] outdoor (0150.jpg).
Figure 12: Image Deraining Results on Rain100L [21] (035.png).
Figure 13: Image Denoising Results on BSD68 [52] (0060.png).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Input InstructIR (Output)
Figure 14: Image Enhancement Results. We provide qualitative samples from the MIT5K Dataset [5].
Instruction: “the image is too dark, it has poor illumination, can you make it brighter?”
Refer to caption Refer to caption
Refer to caption Refer to caption
Instruction: “Please, reduce the motion in this image so it is more clear”
Refer to caption Refer to caption
Figure 15: Additional high-resolution qualitative results using the LOL [83] dataset (low-light image enhancement), and the GoPro [57] dataset (motion deblurring). We provide the corresponding natural language instructions.
[…]
Supplementary Material
We define our loss functions in the paper Sec. 4.1. Our training loss function is $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{ce}$, which includes the loss function of the image model ($\mathcal{L}_{1}$), and the loss function for intent (task/degradation) classification ($\mathcal{L}_{ce}$) given the prompt embedding. We provide the loss evolution plots in Figures 16 and 17. In particular, in Figure 17 we can observe how the intent classification loss (i.e. predicting the task (or degradation) given the prompt), tends to 0 very fast, indicating that our language model component can infer easily the task given the instruction.
Text Discussion
We believe our instructions cover a wide range of expressions (technical, common language, ambiguous, etc). The language model works properly on real-world instructions. Therefore, we believe the language for this specific task is self-constrained, and easier to understand and to model in comparison to other “open” tasks such as image generation.
Model Design
How the 6D variant does Super-Resolution?. We degraded the input images by downsampling and re-upsampling using Bicubic interpolation. Given a LR image, we updample it using Bicubic, then InstructIR can recover some details.
Encoder Deraining Denoising Deblurring LOL
BGE-micro 36.84/0.973 31.40/0.887 29.40/0.886 23.00/0.836
ALL-MINILM 36.82/0.972 31.39/0.887 29.40/0.886 22.98/0.836
CLIP 36.83/0.973 31.39/0.887 29.40/0.886 22.95/0.834
Figure 17: Intent Classification Loss from the instructions. Product of our simple MLP classification head using $\mathbf{e}$. When $\mathcal{L}_{ce}\!\to\!0$ the model uses the learned (optimized) prompt embeddings, and it is optimized mainly based on the image regression loss ($\mathcal{L}_{1}$).
Real-World Generalization.
We evaluate InstructIR as previous works [61, 42, 100]. Also, we find the same limitations as such methods when we process real-world images. Evaluating the model on (multiple) real-world degradations is a future task.
Contemporary Works and Reproducibility.
Additional Visual Results
@@ 7 Conclusion
We present diverse qualitative samples in Figures 1920, and 21. Our method produces high-quality results given images with any of the studied degradations. In most cases the results are better than the reference all-in-one model AirNet [42]. Download all the test results at https://github.com/mv-lab/InstructIR.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Comparison with InstructPix2Pix [4] for instruction-based restoration using the prompt “Remove the noise in this photo”.
@@
Input AirNet [42] PromptIR [61] InstructIR Reference
Figure 21: Dehazing comparisons for all-in-one methods on images from the SOTS outdoor dataset [41].

### v5

@@ 5 Multi-Task Discussion and Study
How does 7D work?
Finally, if we add real image enhancement –a task not related to the previous ones i.e. inverse problems– the performance on the restoration tasks decays slightly. However, the model still achieves state-of-the-art results. Moreover, as we show in Table 5, the performance on this task using the MIT5K [5] dataset is notable, while keeping the performance on the other tasks.
Table 6: Summary ablation study on the multi-task variants of InstructIR that tackle from 3 to 7 tasks.
@@
We compare with task-specific methods for real-world photography enhancement in Table 5, and for real-world low-light image enhancement in Table 7.
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Clean up my image, “Get rid of the grain “Remove the strange
it’s too fuzzy." in my photo" spots on my photo"
Refer to caption Refer to caption Refer to caption Refer to caption
“Retouch this image “Reduce the motion “Please get rid of “Reduce the fog in
and improve colors" in this shot" the raindrops" this landmark"
Figure 5: Adversarial and OOD samples for Instruction-based Restoration. InstructIR understands a wide range of instructions for a given task (first row). Given an adversarial or out-of-distribution instruction (second row), the model does not modify the image notably (i.e. performs the identity) –we did not enforce this during training–.
Table 7: Quantitative comparisons with state-of-the-art methods on the LOL [84] dataset for Real-World Low-light Enhancement. Note that InstructIR-7D is a multi-task method, while the other methods are task-specific. Table based on [82].
@@
Refer to caption Refer to caption Refer to caption
Input (1) “Clear the rain from my picture" $\longrightarrow$ (2) “Make it pop"
Refer to caption Refer to caption Refer to caption
(1) “Retouch it as a photographer" $\longrightarrow$ “Can you remove the raindrops?" $\longrightarrow$ “Increase the resolution"
Refer to caption Refer to caption Refer to caption
Input (1) “My image is too dark, fix it" $\longrightarrow$ (2) “Apply a tonemap"
Figure 6: Control via instructions. We can prompt multiple instructions (in sequence) to restore and enhance the images. This provides additional control. We show two examples of multiple instructions applied to the “Input" image -from left to right-.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Comparison with InstructPix2Pix [4] using the prompt “Remove the noise in this photo". Real-case image from RealSRSet [45].
5.1 On the Effectiveness of Instructions
Thanks to our integration of human instructions, users can control how to enhance the images. We provide examples in Figures 5 and 6, where we show the potential of our method to restore and enhance images in a controllable manner.
This implies an advancement w.r.t classical (deterministic) image restoration methods. Classical deep restoration methods lead to a unique result, thus, they do not allow to control how the image is processed. We also compare InstructIR with InstructPix2Pix [4] (a diffusion-based generative model) in Figure 7.
Qualitative Results.
We provide diverse qualitative results for several tasks, and we compare with all-in-one and task-specific methods. In Figure 10, we show results on the LOL [84] dataset. In Figure 11, we compare methods on the motion deblurring task using the GoPro [58] dataset. In Figures 9 and 12, we compare with different methods for the dehazing task on SOTS (outdoor) [42]. In Figure 13, we compare with image restoration methods for deraining on Rain100L [21]. Finally, we show denoising results in Figure 14. In this qualitative analysis, we use our single InstructIR-5D model to restore all the images.
Refer to caption Refer to caption Refer to caption Refer to caption
“Increase the brightness of my photo please, is it totoro?”  |  “my image is too dark, fix it”
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Real-world samples of image restoration and enhancement using InstructIR.
Input AirNet [43] PromptIR [62] InstructIR Reference
Figure 9: Dehazing comparisons for all-in-one restoration methods on SOTS [42].
Limitations
As with previous all-in-one methods, our model struggles to process images with more than one degradation (i.e. complex real-world images), or unknown out-of-distribution degradations, yet this is a common limitation among the related methods. However, we believe that these limitations can be surpassed with more realistic training data, and scaling the model’s complexity.
6 Conclusion
We present a novel approach that uses natural human-written instructions to guide the image restoration model. Given a prompt, our multi-task model can recover high-quality images from their degraded counterparts, considering multiple degradations. We achieve state-of-the-art results on several restoration tasks, demonstrating the power and flexibility of instruction guidance. Our results represent a new benchmark for text-guided image restoration.
Figure 10: Real Low-light Enhancement Results. LOL [84] testset (748.png). AirNet [43] and IDR [102] are well-known all-in-one restoration methods. NAFNet [9] is equivalent to InstructIR without text conditions (i.e. our image-only backbone).
Figure 11: Image Deblurring Results. GoPro [58] dataset.
Figure 12: Image Dehazing Results. SOTS [42] outdoor dataset (0150.jpg).
Figure 13: Image Deraining Results on Rain100L [21] (035.png).
Figure 14: Image Denoising Results on BSD68 [53] (0060.png).
Acknowledgments
This work was partly supported by the The Humboldt Foundation (AvH). Work done at the University of Würzburg. Marcos Conde is also supported by Sony Interactive Entertainment, FTG.
Appendix 0.A Additional Training Details and Ablations
We define our loss functions in the paper Sec. 4.1. Our training loss function is $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{ce}$, which includes the loss function of the image model ($\mathcal{L}_{1}$), and the loss function for intent (task/degradation) classification ($\mathcal{L}_{ce}$) given the prompt embedding. We provide the loss evolution plots in Figures 15 and 16. In particular, in Figure 16 we can observe how the intent classification loss (i.e. predicting the task (or degradation) given the prompt), tends to 0 very fast, indicating that our language model component can infer easily the task given the instruction.
Text Discussion
We believe our instructions cover a wide range of expressions (technical, common language, ambiguous, etc). The language model works properly on real-world instructions. Therefore, we believe the language for this specific task is self-constrained, and easier to understand and to model in comparison to other “open" tasks such as image generation.
Model Design
Encoder Deraining Denoising Deblurring LOL
BGE-micro 36.84/0.973 31.40/0.887 29.40/0.886 23.00/0.836
ALL-MINILM 36.82/0.972 31.39/0.887 29.40/0.886 22.98/0.836
CLIP 36.83/0.973 31.39/0.887 29.40/0.886 22.95/0.834
Figure 16: Intent Classification Loss from the instructions. Product of our simple MLP classification head using $\mathbf{e}$. When $\mathcal{L}_{ce}\!\to\!0$ the model uses the learned prompt embeddings, and it is optimized mainly using the image regression loss ($\mathcal{L}_{1}$).
How the 6D variant does Super-Resolution?. We degraded the input images by downsampling and re-upsampling using Bicubic interpolation. Given a LR image, we updample it using Bicubic, then InstructIR can recover some details. As we discuss in the paper, adding this task helps the main task of deblurring.
Contemporary Works and Reproducibility.
0.A.1 Additional Ablation Studies
We provide ablation studies and comparison with more task-specific methods in Tables 9 (image denoising) and Table 10 (image deblurring and dehazing).
@@ 0.A.1 Additional Ablation Studies
Appendix 0.B Additional Visual Results
@@
We present diverse qualitative samples in Figures 1718. Our method produces high-quality results given images with any of the studied degradations. In most cases the results are better than the reference all-in-one model AirNet [43], and the recent SOTA PromptIR [62]. Also we compare with InstructPix2Pix [4] (diffusion-based) in Figure 20 using real-world cases. In Figure 19, we test our method on real-world samples for image dehazing.
0.B.1 Efficiency Analysis
We can process FHD images under 1s on consumer-grade GPUs (12-24Gb). We are also notably faster and more efficient than the SOTA method PromptIR [62] with  2x less parameters (16M vs. 35M), and  1.6x less operations.
Table 11: Inference cost comparison. Some numbers are from [9].
Method MPRNet MIRNet Restormer PromptIR NAFNet InstructIR
MACs(G) 588 786 140 160 65 100
@@ 0.B.1 Efficiency Analysis
Refer to caption Refer to caption Refer to caption
Input Hazy RIDCP [85] InstructIR
Refer to caption Refer to caption Refer to caption
Input Hazy PSD [12] InstructIR
Figure 19: Real Image dehazing comparisons. These are real-world samples without ground-truth. Our method achieves pleasant results as generative models such as RIDCP [85] based on VQGAN. Sample from the RTTS dataset [42]. We use the instruction “remove and haze and mist from this photo please”.
Instruction: “Reduce the noise in this photo" – Basic & Precise
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Remove the tiny dots in this image" – Basic & Ambiguous
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Improve the quality of this image" – Real user (Ambiguous)
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “restore this photo, add details" – Real user (Precise)
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Enhance this photo like a photographer" – Basic & Precise
Refer to caption Refer to caption Refer to caption Refer to caption
Input InstructIR (ours)  $S_{I}\!=\!5$  $S_{I}\!=\!7$
Figure 20: Comparison with [4] for instruction-based restoration using the prompt. Real-world samples from the RealSRSet [81, 45]. We use our 7D variant. We run [4] using two configurations where we vary the weight of the image component hoping to improve fidelity: $S_{I}\!=\!5$ and $S_{I}\!=\!7$ (also known as Image CFG), this parameters helps to enforce fidelity and reduce hallucinations.